OpenStudy (anonymous):

what is the solution of the portion 14/12 =d/48

OpenStudy (anonymous):

\[\frac{ 14 }{ 12 }=\frac{ d }{ 48 }\] Cross multiply 12d=14(48) 12d=252 Divide by 12 on both sides d=21

OpenStudy (kirbykirby):

^I think you made a mistake multiplying 14 and 48 which is 672.